Subject: APO 3.5m users committee minutes, Feb 2013
From: Suzanne Hawley
Submitted: Sun, 24 Feb 2013 19:52:29 -0800 (PST)
Message number: 1305
(previous: 1304,
next: 1306
up: Index)
APO 3.5-m Users Committee Phonecon, 2/4/13
Attending: Suzanne Hawley, Mark Klaene, John Bally, Scott Anderson, Nancy
Chanover (for Jon Holtzman), Michael Strauss, Mary Beth Kaiser, Al Harper,
Remy Indebetouw, Bill Ketzeback, Joe Huehnerhoff, and Bruce Gillespie.
Absent: none
**********************************
User feedback and comments from institutional representatives:
o Washington (Scott Anderson) - Scott had nothing new to report.
o Chicago (Al Harper) - Al had nothing new to report.
o New Mexico State (Nancy Chanover) - Nancy remarked about the recent
problems with NIC-FPS data, and Jon Holtzman is looking into it offline.
John Bally added that using NFS>8 causes crashes of the system. Suzanne
asked if users at Chicago and JHU would please check their recent NIC-FPS
data to see if they are OK.
o Princeton (Michael Strauss) - Michael had nothing new to report.
o Colorado (John Bally) - John mentioned that he had problems with logins
on Newton; the passwords were rejected for sftp. He will check and see if
the problem still exists.
o Virginia (Remy Indebetouw) - Remy had nothing new to report.
o Johns Hopkins (Mary Beth Kaiser) - Mary Beth had nothing new to report.
**********************************
Discussion of telescope/instruments report:
Mark said that aside from a few issues, the telescope systems are working
well. There is a problem with one of the tertiary mirror actuators that
is being monitored. Binning DIS images in the Y direction fails; this is
not a new issue but is getting some attention -- if this is or will be a
problem for anyone, they should let us know. After the ICC for Agile was
repaired, it is now working fine. We seem to have recovered from the
SPIcam Cryotiger gas leak. Updating the earlier report on the echelle
given below, we now have a good vacuum and are cooling the instrument.
We may need to replace the echelle ion vacuum pump, but it is unclear what
the root problem of the vacuum failure was. On the NIC-FPS NSF>8
failures, Stephane Beland has determined that the software seems OK, and
we are now checking the hardware. Updates on the status of NIC-FPS NSF>8
problem will be posted on the updates page on the APO 3.5m instruments
website.
Mark Klaene's detailed report for the reporting period follows:
*************************
3.5-m Telescope and Instruments Highlights
11/28/12 through 1/29/13
0) Overview
A lot of weather-associated lost time. Frequent storm systems about every
three days came through in late December through Mid-January.
Unfortunately, despite all the clouds, little moisture was accumulated.
Minimal clear time was lost due to snow and ice on the roof/lightning
arrestors.
1) Telescope
The telescope performance has been nominal this period. There was one
event that caused a pointing shift that appeared to have its root cause in
a M2 actuator that failed to move due to the cold temperatures. Recovery
was compounded by a less-than-perfect M3 actuator that will need servicing
in the future.
2) Instruments --
DIS has been operating nominally. Both red and blue cameras have been
functioning normally with good vacuum and temperature. Binning of the
chips only works in the Y (rows) direction for an even number of windowed
rows. Choosing an odd number for row binning causes a failure.
The Agile ICC died, which was traced to a failure on the motherboard of
the Dell PC. The ICC was rebuilt and moved to a server platform, and
software was upgraded to make the ICC more maintainable. This should
improve the overall maintainability of Agile since its delivery to APO.
During the ICC work, a communication fiber was found to have low
throughput and was replaced as well.
The SPIcam cryotiger developed a gas leak causing an unexpected warm up.
By pumping down and recharging the unit, the instrument was brought back
up the next day. Cause for the leak has not been definitively identified.
TripleSpec operated with no problems. Recall the slit is still locked in
the 1.1" position.
Echelle is not operational at this writing. The LN2 fill system had a
major failure on January 27 causing the CCD to warm. Staff are working on
identifying and fixing the failure as well as pump and cool the camera.
Given the size and complexity of the instrument this takes several days.
The instrument should be ready by the time of the meeting.
NIC-FPS is operating in shared-risk observing mode due to occasional image
corruption issues. It was discovered that NFS greater than 8 would cause
the ICC to fail which may have been caused by the ICC rebuild that was
done when the ICC failed last summer. Stephane Beland is helping to try
to track down the problem.
**********************************
GIFS training in Q2:
Suzanne said that the GSFC team needs more time to get GIFS operable under
TUI, possibly until May. We need the institutional schedulers to
encourage users to ask for observing time adjacent to the training runs.
Also, we need to know how many users are interested in being trained to
use GIFS.
**********************************
Spectroscopy in LSST era:
Suzanne reported that NOAO is hosting this meeting in mid-April; Rene
Walterbos is going to represent us. The ARC 3.5-m telescope has 2
magnitudes of object overlap with LSST in the jointly seen part of the
sky. One possibility for 3.5m spectroscopy contributions would be to
conduct followup observing of the brighter LSST transients in campaign
mode, like we did during the SDSS-II SNe survey. We also need to think
about how the 3.5-m telescope fits into LSST followup programs when TDSS
starts on the Sloan 2.5-m.
===> ACTION: Committee members to talk with their users re LSST followup
programs, and discuss at next Users Committee meeting.
**********************************
Futures update:
Suzanne mentioned that we are in discussion with several interested
institutions that may want to join ARC or lease telescope time. One
thought is to identify past 3.5-m users who are now at other non-ARC
institutions to see if they are interested in either engaging their new
institutions to be part of ARC, or else find a way for them to form a
participation group with other institutions. If any of the Users
Committee members or their users know of past 3.5-m users to contact,
please do so and alert Suzanne and Rene of any possible interest.
**********************************
DIS slitviewer camera design document:
A white paper describin the DIS slitviewer camera design was sent to the
users committee prior to the meeting. Users can obtain a copy from their
representative. Bill gave the committee a brief introduction to the
proposed upgrade to the DIS slitviewer. The upgrade includes a new camera
that has better blue sensitivity, and the optics upgrade affords a larger
FOV. If there are other motivations or improvements that users would like
us to consider for the slitviewer upgrade, let us know. John thought it
might be useful for the slitviewer to have a filter wheel. Remy asked
about the priority of this project, and Mark said that it is relatively
high priority in CIF work, behind the rewrite of the TCC software.
===> ACTION: Users should send comments about the DIS slitviewer upgrade
to Bill, Mark and Suzanne.
**********************************
ACTION ITEMS:
[open from previous meeting]:
===> ACTION: Send new ARC partner suggestions to Suzanne and Rene.
Status: Open, committee members need to keep sending suggestions.
We need the names of institutions (of any size) and a named contact.
===> ACTION: Users committee members should discuss APOGEE fiber feed to
3.5-m proposal with their users and forward feedback, and ideally name(s)
of interested users who would be willing to participate in an ATI proposal
to Suzanne, Mark and Fred Hearty.
Status: Closed for now, reopen after the APOGEE MRI proposal
request for a 2nd instrument is decided.
[new from this meeting]:
===> ACTION: Committee members to talk with their users re LSST followup
programs, and discuss at next Users Committee meeting.
===> ACTION: Committee members to solicit omments about the DIS
slitviewer upgrade.
**********************************
Next meeting: The next Users Committee phonecon will be on Monday 4 March
at 8:30 AM Pacific Time. The agenda and other materials will be sent to
the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1305 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO